Tier1 Operations Report 2015-11-18

From GridPP Wiki
Jump to: navigation, search

RAL Tier1 Operations Report for 18th November 2015

Review of Issues during the week 11th to 18th November 2015.
  • We have seen high usage of both the OPN link to CERN and the "bypass" link (to Tier2s) in the last few days. The OPN link has been generally busier (but not full) over the last couple of months.
  • Over the last couple of weeks there have been some failures of Atlas batch jobs which report a missing heartbeat.
  • We are investigating a problem that has affected some Castor disk servers ability to make name (DNS & NIS) lookups.
Resolved Disk Server Issues
  • Both GDS663 (AtlasTape - D0T1) and GDSS676 (CMSTape - D0T1) were taken out of service on the 11th November. In both cases the servers were not responding correctly to DNS lookups. The servers were returned to service during Wednesday, 11th November. However, the cause of the problem has not been understood.
  • GDSS654 (LHCbRawRDst - D0T1) was also taken out of service on Wednesday 11th November. A second disk was found to have failed while an earlier replaced disk was being rebuilt. The server was returned to service on Friday afternoon, 13th November after a third disk was replaced as it was showing a large number of media errors.
Current operational status and issues
  • There is a problem seen by LHCb of a low but persistent rate of failure when copying the results of batch jobs to Castor. There is also a further problem that sometimes occurs when these (failed) writes are attempted to storage at other sites.
  • The intermittent, low-level, load-related packet loss seen over external connections is still being tracked. Likewise we have been working to understand some remaining low level of packet loss seen within a part of our Tier1 network.
Ongoing Disk Server Issues
  • GDSS707 (AtlasDataDisk - D1T0) has been out of production since Friday (16th Oct). The server was drained and is currently undergoing testing with fabric.
  • GDSS720 (AtlasDataDisk - D1T0) failed on Monday evening (9th Nov). The error looks to be in the CPU. The server has been drained, the CPU swapped, and the server is being re-tested ahead of being returned to service.
Notable Changes made since the last meeting.
  • All Tier1 tape servers are now running Castor 2.1.15.
  • Kernel and errata updates rolled out.
Declared in the GOC DB
  • None
Advanced warning for other interventions
The following items are being discussed and are still to be formally scheduled and announced.
  • Upgrade of remaining Castor disk servers (those in tape-backed service classes) to SL6. This will be transparent to users.
  • Some detailed internal network re-configurations to enable the removal of the old 'core' switch from our network. This includes changing the way the UKLIGHT router connects into the Tier1 network.

Listing by category:

  • Databases:
    • Switch LFC/3D to new Database Infrastructure.
  • Castor:
    • Update SRMs to new version (includes updating to SL6).
    • Update disk servers to SL6 (ongoing)
    • Update to Castor version 2.1.15.
  • Networking:
    • Complete changes needed to remove the old core switch from the Tier1 network.
    • Make routing changes to allow the removal of the UKLight Router.
  • Fabric
    • Firmware updates on remaining EMC disk arrays (Castor, LFC)
Entries in GOC DB starting since the last report.
Service Scheduled? Outage/At Risk Start End Duration Reason
Whole Site SCHEDULED WARNING 18/11/2015 09:30 18/11/2015 10:30 1 hour Warning on site during small change to network configuration.
Open GGUS Tickets (Snapshot during morning of meeting)
GGUS ID Level Urgency State Creation Last Update VO Subject
117601 Green Less Urgent In Progress 2015-11-14 2015-11-16 OPS [Rod Dashboard] Issue detected : org.nordugrid.ARC-CE-SRM-result-/ops/Role=lcgadmin@arc-ce02.gridpp.rl.ac.uk
116866 Green Less Urgent On Hold 2015-10-12 2015-10-19 SNO+ snoplus support at RAL-LCG2 (pilot role)
116864 Yellow Urgent Waiting Reply 2015-10-12 2015-11-12 CMS T1_UK_RAL AAA opening and reading test failing again...
Availability Report

Key: Atlas HC = Atlas HammerCloud (Queue ANALY_RAL_SL6, Template 508); CMS HC = CMS HammerCloud

Day OPS Alice Atlas CMS LHCb Atlas HC CMS HC Comment
11/11/15 100 100 100 100 100 95 100
12/11/15 100 100 100 100 100 97 100
13/11/15 100 100 100 100 100 97 100
14/11/15 100 100 100 100 100 100 n/a
15/11/15 100 100 100 100 100 100 n/a
16/11/15 100 100 100 100 100 100 n/a
17/11/15 100 100 100 100 100 100 100